OpenStudy (anonymous):

Order the group of quadratic functions from widest to narrowest graph. y=-4x^2, y=-3x^2, y=-5x^2 A.Y=-3X^2,Y=-5X^2, Y=-4X^2 B. y=5x^2, y=-4x^2 ,y=-3x^2 C.y=-3x^2,y=-4x^2,y=-5x^2 D.Y=4X^2, y=-3x^2 , y=-5x^2

OpenStudy (anonymous):

(A)

OpenStudy (anonymous):

Check B). You seem to have a typo.

OpenStudy (anonymous):

o

OpenStudy (anonymous):

is that a conformation?

OpenStudy (anonymous):

ok now look lol

OpenStudy (anonymous):

i think B) is y=-5x^2, y=-4x^2, y=-3x^2

OpenStudy (anonymous):

If y=ax^2 the magnitude of a depends the width of the graph........ lower the magnitude of a wider the graph

OpenStudy (anonymous):

so, what do u think?

OpenStudy (anonymous):

@rajathsbhat i dont think that is correct

OpenStudy (anonymous):

it is now lol

OpenStudy (anonymous):

Hey, the length of the latus rectum is 4a, right? So, greater the "a", longer is the latus rectum, wider is the graph!

OpenStudy (anonymous):

now i think it is (B)

OpenStudy (anonymous):

& you are right!

OpenStudy (anonymous):

Look at the sign dude they are negative

OpenStudy (anonymous):

Dude, that only means the parabola in facing down, that's all.

OpenStudy (anonymous):

So, we need to take the magnitude only

OpenStudy (anonymous):

Can u PLZ draw a graph

OpenStudy (anonymous):

for each and see

OpenStudy (anonymous):

I'll try.

OpenStudy (anonymous):

ooo. You r right!, @saurav. @ james, you should write it in the form \[x^{2}=-4ay & then compare "a".

OpenStudy (anonymous):

I thought it was in the form \[y^{2}=-4ax\]

OpenStudy (anonymous):

U forget shakya

OpenStudy (anonymous):

lol. But the answer then is C isn't it?

OpenStudy (anonymous):

YEP...... its C

OpenStudy (anonymous):

The real asker is @85295james. So, do you get it, James?

OpenStudy (anonymous):

James, do you understand why it's C?

OpenStudy (anonymous):

Please reply.

OpenStudy (anonymous):

o thank you